OpenStudy (anonymous):

A total of 273 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (anonymous):

There were 91 adult tickets sold. 273/3 = 91 91x2 = 182 182+91 = 273

OpenStudy (calculusxy):

You can also set up an equation. Let x = the number of adult tickets 2x = the number of student tickets 2x + x = 273

OpenStudy (calculusxy):

Combine the like terms: 2x + x to get 3x Now you have 3x = 273 Do the division property of equality, to isolate x. 273/3 = 3x/3 x=91 tickets

OpenStudy (anonymous):

Why is there so many guests?

OpenStudy (calculusxy):

That was what I was thinking also!

OpenStudy (anonymous):

Lol somebody must be making more tabs

OpenStudy (one098):

It's always someone up to annoy us.

OpenStudy (anonymous):

Yep :)